OpenStudy (anonymous):

Assume a data set is normally distributed with mean 121 and standard deviation 15. If the data set contains 650 data values, approximately how many of the data values will fall within the range 91 to 151?

OpenStudy (anonymous):

\[P(91<X<151)=P\left(\frac{91-121}{15}<\frac{X-121}{15}<\frac{151-121}{15}\right)=P(-2<Z<2)\] This probability gives you the proportion of data values that fall in the range. To get the actual number of data values in the interval, multiply the proportion by the total number of data points.